OpenStudy (anonymous):

how do you find the x intercept of this rational function: R(x) = (x+1)/x(x+1)

OpenStudy (anonymous):

\[R(x) = \frac{\cancel{(x+1)}}{x \cancel{(x+1)}}=\frac{1}{x}\]In genral you find x-intercepts of rational functions by setting the numerator equal to zero. However, in your function, the denominator in lowest terms is a constant, indicating no x-intercepts. See attached graph. Note the domain is is all real numbers except 0 and -1.

OpenStudy (anonymous):

awesome! thank you

OpenStudy (anonymous):

ur welcome :})